# Performance of multiple filter-cavity schemes for frequency-dependent squeezing in gravitational-wave detectors

### v1

@@ III.2 Mode mismatch
Importantly, mode coupling can manifest in two distinct effective ways: so-called quadratic mismatch, due to waist position or size mismatch; or involve higher order aberrations, typically due to astigmatism of the optics, see [34]. The former is modeled using a unitary basis change $\bm{U}$, while the latter usually corresponds to modes which are stopped by baffles and thus can simply be modeled using an absorption coefficient $\Lambda$. In the following, we will focus our discussion on the quadratic type, that is, we assume that input, intra-cavity, and output mismatches are all modeled by unitary matrices $\bm{U}$.
@@
We make the general assumption that mode mismatch occurs when a beam goes from free space into a cavity and vice-versa, because of limited actuation in the mode-matching telescopes and in the possible low-order aberrations that they induce. However, for coupled cavities, the mode mismatch between the two subcavities only depends on the design of the cavity mirrors, which can be made with much lower imperfections and thus this mode mismatch can made negligible (see Appendix D for a quantitative derivation).
@@
We work in several eigenbases when propagating the modes throughout the cavities. We consider the transverse electromagnetic bases of the input field (index $I$), the eigenbasis of the first cavity (index $1$ for 2FC, $c$ for CFC), the eigenbasis of the second cavity (index $2$ for 2FC, $a$ for CFC), the eigenbasis of the output field (index $O$). In order to have tractable insights in our model, we only consider a single higher order mode that encapsulates all the mode mismatch for each basis. We regroup these two modes in a vector $\bm{a}[\Omega]=\begin{bmatrix}a_{1}[\Omega]\\ a_{2}[\Omega]\end{bmatrix}$ where the first element is the fundamental gaussian mode, while the second is the higher order mode, evaluated at the sideband frequency $\Omega$ from the carrier (see Appendix A for details). Mode mismatch is characterized by two parameters, $\Upsilon\in[0,1]$ the amplitude of the mismatch ($\Upsilon=0$ corresponds to no mismatch), and $\psi_{\text{mm}}$ the phase of the mode mismatch. For example, if we are interested in the matching between a free-space beam and a cavity eigenmode, $\psi_{\text{mm}}=0$ means that all the mismatch comes from beam waist size, while $\psi_{\text{mm}}=\pi/2$ means that all the mismatch comes from beam waist longitudinal position [35]. An intermediary value of $\psi_{\text{mm}}$ corresponds to a linear combination of waist size and position mismatch.
@@
Figure 5: Squeezing degradation from quadratic mode mismatch in 2FC (blue, solid envelope) and CFC (red, dashed envelope), for a total length $L_{1}+L_{2}=L_{a}+L_{c}=5\text{km}+5\text{km}=10$ km. The blue and orange regions are determined by calculating squeezing curves for the values of $\Upsilon_{I}$, $\Upsilon_{O}$, $\Upsilon_{12}$ and $\Upsilon_{c}$ from Table 1, and sweeping the mode mismatch phases ($\psi_{mmO}-\psi_{mmI}$) and ($\psi_{mm12}-\psi_{mmI}$). The inter-cavity mode mismatch 2FC is $\Upsilon_{12}=0.01$. Some examples of CFC squeezing curves are plotted in orange. From top to bottom: quantum noise reduction with mismatch; ratio of the lower envelopes of CFC to 2FC; efficiency; dephasing; misphasing.
@@
where $\gamma,\Delta\omega$ are the bandwidth and detuning of the cavity, $\alpha=1-2\Upsilon_{I}+2\beta\sqrt{\Upsilon_{I}\Upsilon_{O}}\,e^{\mathrm{i}\psi_{R}}$ with $\beta=\sqrt{\frac{1-\Upsilon_{I}}{1-\Upsilon_{O}}}\simeq 1$ and $\psi_{R}=\psi_{O}+\psi_{G}-\psi_{I}$. The figures of merit are (see Eqs. 36 and 37)
@@
To study the degradation in 2FC close to the detuning frequencies, notice in the formula Eq. 17, the first cavity “sees” a mismatch of $\bm{U}(\Upsilon_{I},\psi_{mmI})$, while the second cavity sees a mismatch of $\bm{U}(\Upsilon_{12},\psi_{mm12})\bm{U}(\Upsilon_{I},\psi_{mmI})$. This means that, outside of the resonance of the second cavity ($r_{2}\simeq 1$), the response $\bm{F}_{2FC}$ is independent of $\Upsilon_{12}$ and its upper left element can be approximated by Eq. 8. Thus at the resonance of the first cavity, we expect similar degradation than the CFC case (QNR of 5 dB for the values in Table 1). On the contrary, outside of the resonance of the first cavity ($r_{1}\simeq 1$), $F_{2FC}$ can be approximated by $F_{1FC}$, but provided that one makes the substitution $\bm{U}(\Upsilon_{I},\psi_{mmI})\to\bm{U}(\Upsilon_{12},\psi_{mm12})\bm{U}(\Upsilon_{I},\psi_{mmI})$. We then redefine $\alpha=1-2\Upsilon^{\prime}+2\beta^{\prime}\sqrt{\Upsilon^{\prime}\Upsilon_{O}}e^{\mathrm{i}\psi^{\prime}}$ where $\beta^{\prime}=\sqrt{\frac{1-\Upsilon^{\prime}}{1-\Upsilon_{O}}}$ and the expressions of $\Upsilon^{\prime}$ and $\psi^{\prime}$ are given in Eq. 66 of Appendix E. Importantly, the maximum amount of mismatching thus obtained is larger than the sum of the mismatches: $\Upsilon^{\prime}_{\text{max}}=0.088>\Upsilon_{I}+\Upsilon_{12}=0.05$, which means that the degradation at the second cavity’s resonance can be larger than the first’s, as shown in the figure. Numerically, at $\Delta\omega_{2}$, we find, by replacing $\alpha$ with $\alpha^{\prime}$ in the formulas for 1FC, that in the worst case ($\Upsilon^{\prime}=\Upsilon^{\prime}_{\text{max}}$ and $\psi^{\prime}=\pi$), $\eta[\Delta\omega_{2}]=0.74$ and $\Xi[\Delta\omega_{2}]=0.025$, implying that QNR $\sim 3$ dB, as shown in Fig. 5. Importantly, this highlights that mode mismatch does not affect both cavities similarly, as the second cavity may experience larger degradation around its frequencies of resonance. This means that, in order to favor low frequencies, it may be beneficial to place the cavity with the lowest detuning first, opposite to what is shown in Fig. 5.
@@ III.3 Filter cavity length fluctuations
The locking imprecision on the filter cavities induce a length fluctuation $\delta L$ which converts into phase noise for the squeezed field. Using the model developed in Section A.4, we plot its effect on squeezing in Fig. 6. The input squeezing level is 10 dB. We see that overall, cavity length fluctuations of less than 2 pm do not significantly degrade squeezing, neither for CFC nor 2FC, with a marginally better performance of CFC. Importantly, we have assumed that all the cavity lengths ($L_{1}$ and $L_{2}$ for the 2FC; $L_{c}$ and $L_{a}$ for the CFC) have the same fluctuation RMS $\delta L$. In practice, this will depend on the actual control scheme used to stabilize each of these cavities. While we defer the concrete technical control scheme to future work, we note that the range of values chosen for this length noise is realistic, as an FDS cavity controlled down to $\delta L\sim 1$ ppm was achieved in the current generation of GW detectors [9, 8].
Figure 6: Squeezing degradation for length fluctuation $\delta L$ for the 2FC and CFC, for a total length $L_{1}+L_{2}=L_{a}+L_{c}=5\text{km}+5\text{km}=10$ km.
@@ III.4 Full degradation budget comparison
Figure 7: Comparison of CFC and 2FC using the full budget of Table 1.
Having individually studied all three loss contributions, we now determine the full picture of loss degradation in a 2FC and CFC. The response is illustrated in Fig. 7. The non-linear addition of mode mismatches in 2FC, along with the extra optical losses of the FI, ultimately degrade the total squeezing level below that of the CFC scheme. This study thus concludes that, all other parameters and degradation sources being equal, CFC overall performs better than 2FC.
@@ IV Tuned single filter cavity configuration for Einstein Telescope
It has been repeatedly noted in the literature that the practical operation of a GW interferometer in a detuned configuration poses significant technical challenges in the stabilization of the various control loops [36, 37]. In this section, we propose an alternative to ET-LF that corresponds to a tuned configuration, along with a single filter cavity for optimal frequency-dependent squeezing (referred henceforth as the 1FC configuration). This would serve as either 1) an intermediary and easier to control configuration before upgrading to the full detuned setpoint or 2) a fallback solution, should detuned ET-LF prove to be technically difficult. Importantly, we emphasize that the 1FC configuration uses essentially the same infrastructure and excavation footprint as CFC, so that no significant overhead may be required, other than changing the mirror transmissivities. The optical layout is shown in Fig. 8 and compared to CFC: the length of the 1FC cavity is $L_{\text{FC}}=2L=L_{a}+L_{c}$.
@@
Previous studies of tuned ET-LF with 1FC, such as [21, 38], only considered quantum noise degradation due to round trip loss and did not assume equal footprint $L_{\text{1FC}}=L_{a}+L_{c}$ in their comparison to detuned ET-LF + CFC. In the following, we provide a comparison with the full noise budget, which also includes contributions from the interferometer, and we keep the total 1FC/CFC lengths to be equal.
The setpoint and the noise curves for CFC are obtained using the method described in Section III, then converted to redshift units. To maximize the astrophysical reach of tuned ET + 1FC, we optimize on the redshift and use as free parameters the signal recycling transmissivity $T_{\text{SRM}}$, the transmissivity of the 1FC’s input mirror $T_{\text{FC}}$ as well as the detuning of 1FC $\Delta\omega_{\text{FC}}$. These are natural parameters to optimize [19, 39, 36, 28]. The results for total length of $L_{\text{1FC}}=L_{a}+L_{c}=2\text{ km}$ and $10$ km are presented in Fig. 9, while the optimized parameters for 1FC are shown in Table 2. Note that, thanks to the fact that the 1FC experiences less overall quantum degradation, it is possible to push the injected squeezing up to 19 dB without suffering from significant dephasing or phase noise.
Figure 9: Astrophysical reach of ET when ET-LF is tuned or not detuned, assuming the total length of 2 FC and length of 1FC will be identical. SEC refers to ”Signal Extraction Cavity”.
@@
$T_{\mathrm{SRM}}$ Transmissivity of signal recycling mirror $50$
$T_{\text{FC}}^{\mathrm{2km}}$ 2km FC input mirror transmissivity $0.08$
$\Delta\omega_{\text{FC}}^{\mathrm{2km}}$ 2km FC detuning $4.78$ Hz
$T_{\text{FC}}^{\mathrm{5km}}$ 10km FC input mirror transmissivity $0.40$
$\Delta\omega_{\text{FC}}^{\mathrm{10km}}$ 10km FC detuning $4.78$ Hz
$r_{\text{tuned}}$ Injected squeezing $19$ dB
Table 2: Optimized parameters for tuned ET-LF + 1FC.
[…]
As a conclusion, this study reinforces the relevance of aiming for the tuned + 1FC configuration as an alternative or an intermediary for the detuned configuration.

### v2

@@ III.2 Mode mismatch
Importantly, mode coupling can manifest in two distinct effective ways: so-called quadratic mismatch, due to waist position/size mismatch or other large-scale defects of the optics; or involve higher order aberrations, typically due to point defects and high-order scattering, see [35]. The former is modeled using a unitary basis change $\bm{U}$, while the latter usually corresponds to modes which are stopped by baffles and thus can simply be modeled using an absorption coefficient $\Lambda$. In the following, we will focus our discussion on the quadratic type, that is, we assume that input, intra-cavity, and output mismatches are all modeled by unitary matrices $\bm{U}$.
@@
We make the general assumption that mode mismatch occurs when a beam goes from free space into a cavity and vice-versa, because of limited actuation in the mode-matching telescopes and in the possible low-order aberrations that they induce. However, for coupled cavities, the mode mismatch $\Upsilon_{a}$ between the two subcavities only depends on the design of the cavity mirrors, which can be made with much lower imperfections and thus this mode mismatch can made negligible (see Appendix D for a quantitative derivation).
@@
We work in several eigenbases when propagating the modes throughout the cavities. We consider the transverse electromagnetic bases of the input field (index $I$), the eigenbasis of the first cavity (index $1$ for 2FC, $c$ for CFC), the eigenbasis of the second cavity (index $2$ for 2FC, $a$ for CFC), the eigenbasis of the ”output” field which exits the FDS system and gets injected into the interferometer (index $O$). In order to have tractable insights in our model, we only consider a single higher order mode that encapsulates all the mode mismatch for each basis. We regroup these two modes in a vector $\bm{a}[\Omega]=\begin{bmatrix}a_{1}[\Omega]\\ a_{2}[\Omega]\end{bmatrix}$ where the first element is the fundamental gaussian mode, while the second is the higher order mode, evaluated at the sideband frequency $\Omega$ from the carrier (see Appendix A for details). Mode mismatch is characterized by two parameters, $\Upsilon\in[0,1]$ the amplitude of the mismatch ($\Upsilon=0$ corresponds to no mismatch), and $\psi_{\text{mm}}$ the phase of the mode mismatch. For example, if we are interested in the matching between a free-space beam and a cavity eigenmode, $\psi_{\text{mm}}=0$ means that all the mismatch comes from beam waist size, while $\psi_{\text{mm}}=\pi/2$ means that all the mismatch comes from beam waist longitudinal position [36]. An intermediary value of $\psi_{\text{mm}}$ corresponds to a linear combination of waist size and position mismatch.
@@
Figure 5: Squeezing degradation from quadratic mode mismatch in 2FC (blue, solid envelope) and CFC (red, dashed envelope), for a total length $L_{1}+L_{2}=L_{a}+L_{c}=5\text{km}+5\text{km}=10$ km. The blue and orange regions are determined by calculating squeezing curves for the values of $\Upsilon_{I}$, $\Upsilon_{O}$, $\Upsilon_{12}$ and $\Upsilon_{a}$ from Table 1, and sweeping the mode mismatch phases ($\psi_{mmO}-\psi_{mmI}$) and ($\psi_{mm12}-\psi_{mmI}$). The inter-cavity mode mismatch 2FC is $\Upsilon_{12}=0.01$. Some examples of CFC squeezing curves are plotted in orange. From top to bottom: quantum noise reduction with mismatch; ratio of the lower envelopes of CFC to 2FC; efficiency; dephasing; misphasing.
@@
where $\gamma,\Delta\omega$ are the bandwidth and detuning of the cavity, $\alpha=1-2\Upsilon_{I}+2\beta\sqrt{\Upsilon_{I}\Upsilon_{O}}\,e^{\mathrm{i}\psi_{R}}$ with $\beta=\sqrt{\frac{1-\Upsilon_{I}}{1-\Upsilon_{O}}}\simeq 1$ and $\psi_{R}=\psi_{O}+\psi_{G}-\psi_{I}$. The figures of merit are (see Eqs. 35 and 36)
@@
To study the degradation in 2FC close to the detuning frequencies, notice in the formula Eq. 17, the first cavity “sees” a mismatch of $\bm{U}(\Upsilon_{I},\psi_{mmI})$, while the second cavity sees a mismatch of $\bm{U}(\Upsilon_{12},\psi_{mm12})\bm{U}(\Upsilon_{I},\psi_{mmI})$. This means that, outside of the resonance of the second cavity ($r_{2}\simeq 1$), the response $\bm{F}_{2FC}$ is independent of $\Upsilon_{12}$ and its upper left element can be approximated by Eq. 8. Thus at the resonance of the first cavity, we expect similar degradation than the CFC case (QNR of 5 dB for the values in Table 1). On the contrary, outside of the resonance of the first cavity ($r_{1}\simeq 1$), $F_{2FC}$ can be approximated by $F_{1FC}$, but provided that one makes the substitution $\bm{U}(\Upsilon_{I},\psi_{mmI})\to\bm{U}(\Upsilon_{12},\psi_{mm12})\bm{U}(\Upsilon_{I},\psi_{mmI})$. We then redefine $\alpha=1-2\Upsilon^{\prime}+2\beta^{\prime}\sqrt{\Upsilon^{\prime}\Upsilon_{O}}e^{\mathrm{i}\psi^{\prime}}$ where $\beta^{\prime}=\sqrt{\frac{1-\Upsilon^{\prime}}{1-\Upsilon_{O}}}$ and the expressions of $\Upsilon^{\prime}$ and $\psi^{\prime}$ are given in Eq. 77 of Appendix E. Importantly, the maximum amount of mismatching thus obtained is larger than the sum of the mismatches: $\Upsilon^{\prime}_{\text{max}}=0.088>\Upsilon_{I}+\Upsilon_{12}=0.05$, which means that the degradation at the second cavity’s resonance can be larger than the first’s, as shown in the figure. Numerically, at $\Delta\omega_{2}$, we find, by replacing $\alpha$ with $\alpha^{\prime}$ in the formulas for 1FC, that in the worst case ($\Upsilon^{\prime}=\Upsilon^{\prime}_{\text{max}}$ and $\psi^{\prime}=\pi$), $\eta[\Delta\omega_{2}]=0.74$ and $\Xi[\Delta\omega_{2}]=0.025$, implying that QNR $\sim 3$ dB, as shown in Fig. 5. This highlights that mode mismatch does not affect both cavities similarly, as the second cavity may experience larger degradation around its frequencies of resonance. This means that, in order to favor low frequencies, it may be beneficial to place the cavity with the lowest detuning first, opposite to what is shown in Fig. 5.
@@ III.3 Filter cavity length fluctuations
The locking imprecision on the filter cavities induce a length fluctuation $\delta L$ which converts into phase noise for the squeezed field. Using the model developed in Section A.4, we plot its effect on squeezing in Fig. 6. The input squeezing level is 10 dB. We see that overall, cavity length fluctuations of less than 2 pm do not significantly degrade squeezing, neither for CFC nor 2FC, with a marginally better performance of CFC. Importantly, we have assumed that all the cavity lengths ($L_{1}$ and $L_{2}$ for the 2FC; $L_{c}$ and $L_{a}$ for the CFC) have the same fluctuation RMS $\delta L$. In practice, this will depend on the actual control scheme used to stabilize each of these cavities. While we defer the concrete technical control scheme to future work, we note that the range of values chosen for this length noise is realistic, as an FDS cavity controlled down to $\delta L\sim 1$ pm was achieved in the current generation of GW detectors [11, 7].
Figure 6: Squeezing degradation for length fluctuation $\delta L$ for 2FC (solid lines) and CFC (dashed lines), for a total length $L_{1}+L_{2}=L_{a}+L_{c}=5\text{km}+5\text{km}=10$ km.
@@ III.4 Full degradation budget comparison
Figure 7: Comparison of quantum noise degradation in each ET-LF interferometer, between CFC (red dashed curve) and 2FC (blue solid curve), using the full budget of Table 1. We have chosen the arms angle to be 90o for simplicity, so that the strain is equal to the differential arm length.
Having individually studied all three loss contributions, we now determine the full picture of loss degradation in a 2FC and CFC. The response is illustrated in Fig. 7. The non-linear addition of mode mismatches in 2FC, along with the extra optical losses of the Injection Faraday isolator, ultimately degrade the total squeezing level below that of the CFC scheme. We also include interferometer losses present due to the injection Faraday isolators (see Fig. 1)  [^1] We retain a Faraday injection loss of $1\%$ while in the rest of the optical layout, each single-pass individual Faraday isolator incurs a loss of $0.5\%$. This is consistent because in practice, to greatly minimize backreflection, we actually require two Faraday isolators in the injection stage, so $2\times 0.5=1\%$., in the arm cavities, in the Signal Extraction Cavity and at the readout (see Table 1 for their values and Appendix F for explicit derivation of the loss contributions). This study thus concludes that, all other parameters and degradation sources being equal, CFC overall performs similarly or slightly better than 2FC.
@@ IV Tuned single filter cavity configuration for Einstein Telescope
It has been repeatedly noted in the literature that the practical operation of a GW interferometer with a detuned signal extraction cavity poses significant technical challenges in the stabilization of the various control loops [38, 39]. In this section, we propose an alternative to ET-LF that corresponds to a tuned configuration, along with a single filter cavity for optimal frequency-dependent squeezing (referred henceforth as the 1FC configuration). This would serve as either 1) an intermediary and easier-to-control configuration before upgrading to the full detuned setpoint or 2) a fallback solution, in case the detuned setpoint of ET-LF is shown to be technically difficult to stabilize. Importantly, we note that the 1FC configuration uses essentially the same infrastructure and excavation footprint as CFC, so that no significant overhead may be required in terms of hardware, other than changing the mirror transmissivities. The optical layout is shown in Fig. 8 and compared to CFC: the length of the 1FC cavity is $L_{\text{1FC}}=2L=L_{a}+L_{c}$.
@@
Previous studies of tuned ET-LF with 1FC, such as [23, 40], only considered quantum noise degradation due to round trip loss and did not assume equal footprint $L_{\text{1FC}}=L_{a}+L_{c}$ in their comparison to detuned ET-LF + CFC/2FC. By keeping the total 1FC/CFC/2FC lengths to be equal and also including dephasing loss induced by mode mismatch (same values as in Table 1), we find that an increase in injected squeezing level negatively impacts the total sensitivity, as the mode mismatch phases are in general unconstrained, as shown in Fig. 9, where an injected squeezing level of 20~20 dB (as initially studied in [24]) may lead to worse performance for some choices of mismatch phases than lower levels (12 dB). This confirms that, even in the case of a single filter cavity and tuned ET-LF, mode mismatch remains a limitation for squeezing injection.
Figure 9: Quantum noise of ET-LF when the signal extraction cavity is tuned and a single 10 km filter cavity is used for frequency dependent squeezing. Different color curves show quantum noise with different level of input squeezing. The shaded area represents the variation of quantum noise due to the unconstrained mode mismatch phases. All parameters except the tuning of the signal extraction cavity $\phi_{\text{SEC}}$ are identical to Table 1.
We can further provide a comparison with the full noise budget, which includes contributions from quantum noise and other sources related to the interferometer such as residual gas noise, thermal noise, newtonian noise, seismic noise. The quantum noise curves for CFC and 2FC are obtained using the method described in Section III, then added to the other classical noises (see [40]) and converted to redshift units. To maximize the astrophysical reach of tuned ET + 1FC, we optimize on the redshift and use as free parameters the signal recycling transmissivity $T_{\text{SRM}}$, the transmissivity of the 1FC’s input mirror $T_{\text{FC}}$, the injected squeezing level $r_{\text{tuned}}$ as well as the detuning of 1FC $\Delta\omega_{\text{FC}}$. These are well-studied parameters to optimize over in previous and current generation interferometric detectors [17, 41, 38, 31]. We use a particle swarm optimization method [42] to maximize the astrophysical reach for total masses between 10 and 30 solar masses. We also take into account the redshift contribution of the High-Frequency Einstein Telescope interferometer, as well as the triangular geometry of the full Einstein Telescope layout – which simply amounts to an overall scaling factor [43]. The results for total length of $L_{\text{1FC}}=L_{a}+L_{c}=L_{1}+L_{2}=10$ km are presented in Fig. 10, while the optimized parameters for 1FC are shown in Table 2  [^2] We recall that in this study, we consider a single filter cavity for frequency dependent squeezing; the optimized value of $T_{\text{SRM}}$ could be larger had we considered two filter cavities, since in the limit $T_{\text{SRM}}\to 1$ we obtain a power-recycled Michelson interferometer whose optimal squeezing response can only be obtained with two filter cavities [6]..
We see from Fig. 10 that the CFC and 2FC both perform better (up to a factor of 2 in redshift) in the middle-mass range $10M_{\odot},\ 30M_{\odot}]$ due to their enhanced sensitivity in the corresponding frequency range of these mergers, while the tuned configuration offers comparable or better performance outside of this range. Nevertheless, we note that the values of mode mismatch have been taken to be conservative in this study, such that if these values are improved in practice, better performance for both tuned and detuned configurations may be expected. Furthermore, the tuned configuration still achieves orders-of-magnitude redshift improvement over current generation detectors such as Advanced LIGO (maximum redshift of the order of 1), and goes beyond the star formation epoch ($\sim 20$ redshift), which confirms this configuration’s scientific interest.
Figure 10: Astrophysical reach of ET when ET-LF is tuned or not detuned, assuming identical total length of CFC, 2FC and 1FC (10 km) and a geometry given by 60o interferometers in a triangular network. The shaded areas represent the possible redshifts due to the unconstrained mode mismatch angles for each of the configurations.
@@
$T_{\text{FC}}$ FC input mirror transmissivity $0.37$ %
$\Delta\omega_{\text{FC}}^{\mathrm{10km}}$ FC detuning $4.20$ Hz
$r_{\text{tuned}}$ Injected squeezing $12$ dB
$T_{\text{SRM}}$ SR mirror transmissivity $44$ %
Table 2: Optimized parameters for tuned ET-LF + 1FC. Apart from these values and the SEC detuning (set here to $\phi_{\text{SEC}}=\pi/2$ rad), the other interferometer parameters are identical to Table 1.